### v3

@@ 5. Dominant sets and oriented exchange graphs
In this section, we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$, and let $\mathcal{A}$ be the cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. We fix a rooted vertex $t_{0}$ of $\mathbb{T}_{n}$, that is, we fix an initial seed $({\bf x}_{t_{0}},\widetilde{B}_{t_{0}})$ of $\mathcal{A}$. We will first introduce the dominant sets for seeds of cluster algebras as an replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory and as an replacement of inversion sets in the study of right weak order on Weyl groups. Then we prove the main results on oriented exchange graphs.
@@ 5.1. Oriented exchange graphs are acyclic
Two seeds $({\bf x}_{t},\widetilde{B}_{t})$ and $({\bf x}_{w},\widetilde{B}_{w})$ of $\mathcal{A}$ are said to be equivalent if they are the same up to a relabeling. We denote by $[{\bf x}_{t},\widetilde{B}_{t}]$ the equivalent class of $({\bf x}_{t},\widetilde{B}_{t})$.
@@
For a seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{A}$, we denote by $C_{t}=C_{t}^{t_{0}}$ the $C$-matrix of the vertex $t$ with respect to the rooted vertex $t_{0}$.
Definition 5.1 (Green mutation and exchange quiver).
(i) A seed mutation $\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ in $\mathcal{A}$ is called a green mutation, if the $k$th column of the $C$-matrix $C_{t}=C_{t}^{t_{0}}$ is a non-negative vector. Otherwise, it is called a red mutation.
(ii) The exchange quiver (or oriented exchange graph) of $\mathcal{A}$ with initial seed $({\bf x}_{t_{0}},\widetilde{B}_{t_{0}})$ is the quiver $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose arrows correspond to green mutations.
We remark that if two seeds $({\bf x}_{t},\widetilde{B}_{t})$ and $({\bf x}_{w},\widetilde{B}_{w})$ are related by once mutation, then either the mutation from $({\bf x}_{t},\widetilde{B}_{t})$ to $({\bf x}_{w},\widetilde{B}_{w})$ is a green mutation or the mutation from $({\bf x}_{w},\widetilde{B}_{w})$ to $({\bf x}_{t},\widetilde{B}_{t})$ is a green mutation.
For a vector ${\bf h}=(h_{1},\ldots,h_{m})^{T}\in\mathbb{Z}^{m}$, we denote by ${\bf h}^{\circ}:=(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}^{n}$ the principal part of ${\bf h}$.
Denote by $\widetilde{\mathcal{X}}$ the set of cluster variables (frozen and unfrozen) of $\mathcal{A}$ and by $\mathcal{X}$ the set of unfrozen cluster variables of $\mathcal{A}$. Given a cluster monomial $u={\bf x}_{t}^{\bf h}$ in ${\bf x}_{t}$, that is, ${\bf h}^{\circ}=(h_{1},\ldots,h_{n})\in\mathbb{Z}_{\geq 0}^{n}$, we define two subsets of $\mathcal{X}$.
  • The subset $\operatorname{\mathsf{supp}}(u)=\{x_{i;t}\mid i\in[1,n],\;h_{i}\neq 0\}\subseteq\mathcal{X}$ is called the support of $u={\bf x}_{t}^{\bf h}$. We remark that this set only depends on $u$, not on the choice of $t\in\mathbb{T}_{n}$ and ${\bf h}\in\mathbb{Z}^{m}$ such that $u={\bf x}_{t}^{\bf h}$.
  • For a vertex $w\in\mathbb{T}_{n}$ and a cluster variable $z$, recall that
    $(z\mid\mid u)_{F}=F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{z}^{w}].$
    The subset $\operatorname{\mathsf{dom}}^{w}(u)=\{z\in\mathcal{X}\mid F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0\}\subseteq\mathcal{X}$ is called the dominant set of $u={\bf x}_{t}^{\bf h}$ with respect to vertex $w\in\mathbb{T}_{n}$.
@@
Let $({\bf x}_{t},\widetilde{B}_{t})$ be a seed of $\mathcal{A}$. The dominant set $\operatorname{\mathsf{dom}}^{w}[t]$ of $({\bf x}_{t},\widetilde{B}_{t})$ with respect to a vertex $w\in\mathbb{T}_{n}$ is defined to be the dominant set of the multiplicity free cluster monomial $u_{t}=\prod_{i=1}^{n}x_{i;t}$ supported on the unfrozen part of ${\bf x}_{t}$, i.e.,
$\operatorname{\mathsf{dom}}^{w}[t]:=\operatorname{\mathsf{dom}}^{w}(u_{t})=\{z\in\mathcal{X}\mid F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u_{t}}^{w}]=0\}.$
Clearly, if $[{\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}}]=[{\bf x}_{t},\widetilde{B}_{t}]$, then $\operatorname{\mathsf{dom}}^{w}[t^{\prime}]=\operatorname{\mathsf{dom}}^{w}[t]$. Recall that for a cluster ${\bf x}_{t}$, we denote by ${\bf x}_{t}^{\circ}=(x_{1;t},\ldots,x_{n;t})$ the unfrozen part of ${\bf x}_{t}=(x_{1;t},\ldots,x_{m;t})$. If we want to view ${\bf x}_{t},{\bf x}_{t}^{\circ}$ as sets, we will write
$[{\bf x}_{t}]:=\{x_{i;t}\mid i\in[1,m]\}\;\;\;\text{and}\;\;\;[{\bf x}_{t}^{\circ}]:=\{x_{j;t}\mid j\in[1,n]\}.$
@@
Let $u={\bf x}_{t}^{\bf h}$ be a cluster monomial of $\mathcal{A}$. Then the following statements hold for any vertex $w\in\mathbb{T}_{n}$.
@@
    $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$ and in particular, $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}[t]$.
@@
    $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$ and in particular, $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}[t]$.
@@
(i) Since $F_{x_{i;w}}^{w}=1$ for $i\in[1,n]$, we have $F_{x_{i;w}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0$. Thus $x_{i;w}\in\operatorname{\mathsf{dom}}^{w}(u)$ for $i\in[1,n]$, that is, $[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we can get
$[{\bf x}_{w}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u_{t})=\operatorname{\mathsf{dom}}^{w}[t].$
@@
$(x_{i;t}\mid\mid u)_{F}=F_{x_{i;t}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{x_{i;t}}^{w}]$
and each term in this sum is non-negative. Thus we obtain $F_{x_{i;t}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0$, i.e., $x_{i;t}\in\operatorname{\mathsf{dom}}^{w}(u)$. Since $i$ is an arbitrary integer in $[1,n]$, we get $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we can get $[{\bf x}_{t}^{\circ}]\subseteq\operatorname{\mathsf{dom}}^{w}(u_{t})=\operatorname{\mathsf{dom}}^{w}[t]$.
(iii) By definition, an unfrozen cluster variable $z\in\mathcal{X}$ belongs to $\operatorname{\mathsf{dom}}^{w}(u)$ if and only if
$F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0.$
@@
$F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{{\bf x}_{t}^{\bf h}}^{w}]=\sum_{j=1}^{n}h_{j}\cdot F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}].$
Since $h_{j}\geq 0$ and $F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]\geq 0$ for any $j\in[1,n]$, we know that $F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]=0$ if and only if $F_{z}^{w}[(S\mid{\bf 0}){\bf g}_{x_{j;t}}^{w}]=0$ for any $j\in[1,n]$ with $h_{j}\neq 0$, i.e., for any $x_{j;t}\in\operatorname{\mathsf{supp}}(u)$. This is equivalent to say that $z$ belongs to $\bigcap_{x\in\operatorname{\mathsf{supp}}(u)}\operatorname{\mathsf{dom}}^{w}(x)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we see that $\operatorname{\mathsf{dom}}^{w}[t]=\operatorname{\mathsf{dom}}^{w}(u_{t})=\bigcap_{i=1}^{n}\operatorname{\mathsf{dom}}^{w}(x_{i;t})$. ∎
@@
Let $u$ and $u^{\prime}$ be two cluster monomials of $\mathcal{A}$ such that $\operatorname{\mathsf{supp}}(u)=\operatorname{\mathsf{supp}}(u^{\prime})$. Then $\operatorname{\mathsf{dom}}^{w}(u)=\operatorname{\mathsf{dom}}^{w}(u^{\prime})$ for any vertex $w\in\mathbb{T}_{n}$. In particular, $\operatorname{\mathsf{dom}}^{w}[t]=\operatorname{\mathsf{dom}}^{w}(u)$ for any cluster monomial $u$ with $\operatorname{\mathsf{supp}}(u)=[{\bf x}_{t}^{\circ}]$.
@@
For a seed $({\bf x}_{t},\widetilde{B}_{t})$ of $\mathcal{A}$, we denote by $G_{t}=G_{t}^{t_{0}}$ the $G$-matrix of the vertex $t$ with respect to the rooted vertex $t_{0}$ and by
$\mathbb{R}_{\geq 0}G_{t}:=\mathbb{R}_{\geq 0}{\bf g}_{1;t}+\ldots+\mathbb{R}_{\geq 0}{\bf g}_{n;t}\subseteq\mathbb{R}^{n}$
the cone in $\mathbb{R}^{n}$ spanned by the columns of $G_{t}=({\bf g}_{1;t},\ldots,{\bf g}_{n;t})$. We remark that ${\bf g}_{i;t}\in\mathbb{Z}^{n}$ is the principal part of ${\bf g}_{x_{i;t}}^{t_{0}}=\deg^{t_{0}}(x_{i;t})\in\mathbb{Z}^{m}$, i.e., ${\bf g}_{i;t}=({\bf g}_{x_{i;t}}^{t_{0}})^{\circ}$.
We also denote by
$\mathbb{Q}_{>0}G_{t}$ ::=:= $\mathbb{Q}_{>0}{\bf g}_{1;t}+\ldots+\mathbb{Q}_{>0}{\bf g}_{n;t}\subseteq\mathbb{Q}^{n},$
$\mathbb{Z}_{>0}G_{t}$ ::=:= $\mathbb{Z}_{>0}{\bf g}_{1;t}+\ldots+\mathbb{Z}_{>0}{\bf g}_{n;t}\subseteq\mathbb{Z}^{n}.$
@@
Let $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ be a green mutation in $\mathcal{A}$. Then there exists ${\bf v},{\bf v}^{\prime}\in\mathbb{Z}_{>0}^{n}$ such that $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$.
@@
Without loss of generality, we can assume that $k=1$. We write $G_{t}=({\bf g}_{1},{\bf g}_{2},\ldots,{\bf g}_{n})$, where ${\bf g}_{i}\in\mathbb{Z}^{n}$. Since $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})=\mu_{1}({\bf x}_{t},\widetilde{B}_{t})$, the $G$-matrix $G_{t^{\prime}}$ has the form $G_{t^{\prime}}=({\bf g}_{1}^{\prime},{\bf g}_{2},\ldots,{\bf g}_{n})$.
By Theorem 3.10 (i), we know that $\mathbb{R}_{\geq 0}G_{t}$ and $\mathbb{R}_{\geq 0}G_{t^{\prime}}$ are two cones of dimension $n$ in $\mathbb{R}^{n}$ and they have a common facet $\mathcal{C}$ (cone of dimension $n-1$) given by $\mathcal{C}=\mathbb{R}_{\geq 0}{\bf g}_{2}+\ldots+\mathbb{R}_{\geq 0}{\bf g}_{n}$.
Let $S=diag(s_{1},\ldots,s_{n})$ be the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, i.e., $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$. We know that $S$ is a skew-symmetrizer of $B_{t_{0}}$. Let ${\bf c}_{1;t}$ be the first column of the $C$-matrix $C_{t}=C_{t}^{t_{0}}$.
By Theorem 3.10 (iii), we have $SC_{t}S^{-1}(G_{t})^{\rm T}=I_{n}$, equivalently, $(G_{t})^{\rm T}(SC_{t}S^{-1})=I_{n}$. So the first column vector $s_{1}^{-1}(S{\bf c}_{1;t})$ of $SC_{t}S^{-1}$ is an inner normal vector of the facet $\mathcal{C}$ of the cone $\mathbb{R}_{\geq 0}G_{t}$. So when we go along the direction $s_{1}^{-1}(S{\bf c}_{1;t})$, we can cross the common facet $\mathcal{C}$ from the cone $\mathbb{R}_{\geq 0}G_{t^{\prime}}$ to the cone $\mathbb{R}_{\geq 0}G_{t}$.
Now we take a vector ${\bf h}^{\prime}\in\mathbb{Z}_{>0}G_{t^{\prime}}\subseteq\mathbb{R}_{\geq 0}G_{t^{\prime}}$, there exists $\lambda\in\mathbb{Q}_{>0}$ such that
${\bf h}:={\bf h}^{\prime}+\lambda(s_{1}^{-1}(S{\bf c}_{1;t}))\in\mathbb{Q}_{>0}G_{t}\subseteq\mathbb{R}_{\geq 0}G_{t}.$
We can take a positive integer $p$ large enough such that ${\bf a}:=p{\bf h}\in\mathbb{Z}_{>0}G_{t}$ and set ${\bf a}^{\prime}=p{\bf h}^{\prime}\in\mathbb{Z}_{>0}G_{t^{\prime}}$. Then we have $p\lambda s_{1}^{-1}(S{\bf c}_{1;t})={\bf a}-{\bf a}^{\prime}\in\mathbb{Z}^{n}$. Since $({\bf x}_{t^{\prime}},\widetilde{B}_{t^{\prime}})=\mu_{k}({\bf x}_{t},\widetilde{B}_{t})$ is a green mutation, we have ${\bf c}_{1;t}\in\mathbb{Z}_{\geq 0}^{n}$. So we have ${\bf a}-{\bf a}^{\prime}=p\lambda s_{1}^{-1}(S{\bf c}_{1;t})\in\mathbb{Z}_{\geq 0}^{n}$.
Since ${\bf a}\in\mathbb{Z}_{>0}G_{t}$ and ${\bf a}^{\prime}\in\mathbb{Z}_{>0}G_{t^{\prime}}$, there exist ${\bf v},{\bf v}^{\prime}\in\mathbb{Z}_{>0}^{n}$ such that ${\bf a}=G_{t}{\bf v},{\bf a}^{\prime}=G_{t^{\prime}}{\bf v}^{\prime}$. Thus $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}={\bf a}-{\bf a}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$. This completes the proof. ∎

### v4

@@ 7. Dominant sets and oriented exchange graphs
In this section, we introduce the dominant sets for seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory. Then we give applications to the exchange graphs of cluster algebras.
Throughout this section, we will focus on cluster variables and cluster monomials. For this reason, we simply work on cluster algebra with trivial coefficients, rather than $\mathsf{\Lambda}$-cluster algebras. By Proposition 4.30, the $F$-invariant between cluster monomials is well-defined in this case.
@@ 7.1. Oriented exchange graphs are acyclic
Let $B$ be an $n\times n$ skew-symmetrizable matrix with a fixed skew-symmetrizer $S=diag(s_{1},\ldots,s_{n})$. Let $\mathcal{A}$ be the cluster algebra with trivial coefficients whose initial seed is $({\bf x},B)$. The seed of $\mathcal{A}$ at vertex $t\in\mathbb{T}_{n}$ is denoted by $({\bf x}_{t},B_{t})$. In particular, we have $({\bf x}_{t_{0}},B_{t_{0}})=({\bf x},B)$.
Two seeds $({\bf x}_{t},B_{t})$ and $({\bf x}_{w},B_{w})$ of $\mathcal{A}$ are said to be equivalent if they are the same up to relabeling. We denote by ${\bf x}_{t},B_{t}]$ the equivalent class of $({\bf x}_{t},B_{t})$.
@@
For a seed $({\bf x}_{t},B_{t})$ of $\mathcal{A}$, we denote by $C_{t}=C_{t}^{t_{0}}$ the $C$-matrix of the vertex $t$ with respect to the rooted vertex $t_{0}$.
Definition 7.1 (Green mutation and oriented exchange graph).
(i) A seed mutation $\mu_{k}({\bf x}_{t},B_{t})$ in $\mathcal{A}$ is called a green mutation, if the $k$th column of the $C$-matrix $C_{t}=C_{t}^{t_{0}}$ is a non-negative vector. Otherwise, it is called a red mutation.
(ii) The oriented exchange graph of $\mathcal{A}$ with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$ is the quiver $\overrightarrow{\mathcal{H}}^{t_{0}}=\overrightarrow{\mathcal{H}}^{t_{0}}(\mathcal{A})$ whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose arrows correspond to green mutations.
We remark that if two seeds $({\bf x}_{t},B_{t})$ and $({\bf x}_{w},B_{w})$ are related by once mutation, then either the mutation from $({\bf x}_{t},B_{t})$ to $({\bf x}_{w},B_{w})$ is a green mutation or the mutation from $({\bf x}_{w},B_{w})$ to $({\bf x}_{t},B_{t})$ is a green mutation.
Denote by $\mathcal{X}$ the set of cluster variables of $\mathcal{A}$. Given a cluster monomial $u={\bf x}_{t}^{\bf h}$ in ${\bf x}_{t}$, that is, ${\bf h}=(h_{1},\ldots,h_{n})^{T}\in\mathbb{Z}_{\geq 0}^{n}$, we define two subsets of $\mathcal{X}$:
  • The subset $\operatorname{\mathsf{supp}}(u)=\{x_{i;t}\mid i\in[1,n],\;h_{i}\neq 0\}\subseteq\mathcal{X}$ is called the support of $u={\bf x}_{t}^{\bf h}$. We remark that this set only depends on $u$, not on the choice of $t\in\mathbb{T}_{n}$ and ${\bf h}\in\mathbb{Z}^{n}$ such that $u={\bf x}_{t}^{\bf h}$.
  • For a vertex $w\in\mathbb{T}_{n}$ and a cluster variable $z$, recall from Proposition 4.30 that
    $(z\mid\mid u)_{F}=F_{z}^{w}[S{\bf g}_{u}^{w}]+F_{u}^{w}[S{\bf g}_{z}^{w}].$
    The subset $\operatorname{\mathsf{dom}}^{w}(u)=\{z\in\mathcal{X}\mid F_{z}^{w}[S{\bf g}_{u}^{w}]=0\}\subseteq\mathcal{X}$ is called the dominant set of $u={\bf x}_{t}^{\bf h}$ with respect to vertex $w\in\mathbb{T}_{n}$.
Remark 7.2.
Notice that $F_{z}^{w}[S{\bf g}_{u}^{w}]$ is the partial $F$-invariant of the pair $(z,u)$ at vertex $w\in\mathbb{T}_{n}$. That is, the dominant set is defined using the vanishing condition of the partial $F$-invariant.
@@
Let $({\bf x}_{t},B_{t})$ be a seed of $\mathcal{A}$. The dominant set $\operatorname{\mathsf{dom}}^{w}[t]$ of $({\bf x}_{t},B_{t})$ with respect to a vertex $w\in\mathbb{T}_{n}$ is defined to be the dominant set of the multiplicity free cluster monomial $u_{t}=\prod_{i=1}^{n}x_{i;t}$ with full support, i.e.,
$\operatorname{\mathsf{dom}}^{w}[t]\coloneqq\operatorname{\mathsf{dom}}^{w}(u_{t})=\{z\in\mathcal{X}\mid F_{z}^{w}[S{\bf g}_{u_{t}}^{w}]=0\}.$
Clearly, if ${\bf x}_{t^{\prime}},B_{t^{\prime}}]=[{\bf x}_{t},B_{t}]$, i.e., the two seeds are equivalent, then $\operatorname{\mathsf{dom}}^{w}[t^{\prime}]=\operatorname{\mathsf{dom}}^{w}[t]$. If we want to view a cluster ${\bf x}_{t}$ as a set, we will write
${\bf x}_{t}]\coloneqq\{x_{1;t},\ldots,x_{n;t}\}.$
@@
Let $u={\bf x}_{t}^{\bf h}$ be a cluster monomial of $\mathcal{A}$ in seed $({\bf x}_{t},B_{t})$. Then the following statements hold for any vertex $w\in\mathbb{T}_{n}$.
@@
    ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$ and in particular, ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{w}[t]$.
@@
    ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$ and in particular, ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{w}[t]$.
@@
(i) Since $F_{x_{i;w}}^{w}=1$ for $i\in[1,n]$, we have $F_{x_{i;w}}^{w}[S{\bf g}_{u}^{w}]=0$. Thus $x_{i;w}\in\operatorname{\mathsf{dom}}^{w}(u)$ for $i\in[1,n]$, that is, ${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we obtain
${\bf x}_{w}]\subseteq\operatorname{\mathsf{dom}}^{w}(u_{t})=\operatorname{\mathsf{dom}}^{w}[t].$
@@
$(x_{i;t}\mid\mid u)_{F}=F_{x_{i;t}}^{w}[S{\bf g}_{u}^{w}]+F_{u}^{w}[S{\bf g}_{x_{i;t}}^{w}]$
and each term in this sum is non-negative. Thus we obtain $F_{x_{i;t}}^{w}[S{\bf g}_{u}^{w}]=0$, i.e., $x_{i;t}\in\operatorname{\mathsf{dom}}^{w}(u)$. Since $i\in[1,n]$ is arbitrary, we get ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{w}(u)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we obtain ${\bf x}_{t}]\subseteq\operatorname{\mathsf{dom}}^{w}(u_{t})=\operatorname{\mathsf{dom}}^{w}[t]$.
(iii) By definition, a cluster variable $z$ belongs to $\operatorname{\mathsf{dom}}^{w}(u)$ if and only if
$F_{z}^{w}[S{\bf g}_{u}^{w}]=0.$
@@
$F_{z}^{w}[S{\bf g}_{u}^{w}]=F_{z}^{w}[S{\bf g}_{{\bf x}_{t}^{\bf h}}^{w}]=\sum_{j=1}^{n}h_{j}\cdot F_{z}^{w}[S{\bf g}_{x_{j;t}}^{w}].$
Since $h_{j}\geq 0$ and $F_{z}^{w}[S{\bf g}_{x_{j;t}}^{w}]\geq 0$ for any $j\in[1,n]$, we know that $F_{z}^{w}[S{\bf g}_{u}^{w}]=0$ if and only if $F_{z}^{w}[S{\bf g}_{x_{j;t}}^{w}]=0$ for any $j\in[1,n]$ with $h_{j}\neq 0$, i.e., for any $x_{j;t}\in\operatorname{\mathsf{supp}}(u)$. This is equivalent to say that $z$ belongs to $\bigcap_{x\in\operatorname{\mathsf{supp}}(u)}\operatorname{\mathsf{dom}}^{w}(x)$. Hence, $\operatorname{\mathsf{dom}}^{w}(u)=\bigcap_{x\in\operatorname{\mathsf{supp}}(u)}\operatorname{\mathsf{dom}}^{w}(x)$. In particular, if we take $u=u_{t}=\prod_{i=1}^{n}x_{i;t}$, we obtain $\operatorname{\mathsf{dom}}^{w}[t]=\operatorname{\mathsf{dom}}^{w}(u_{t})=\bigcap_{i=1}^{n}\operatorname{\mathsf{dom}}^{w}(x_{i;t})$. ∎
@@
Let $u$ and $u^{\prime}$ be two cluster monomials of $\mathcal{A}$ such that $\operatorname{\mathsf{supp}}(u)=\operatorname{\mathsf{supp}}(u^{\prime})$. Then $\operatorname{\mathsf{dom}}^{w}(u)=\operatorname{\mathsf{dom}}^{w}(u^{\prime})$ for any vertex $w\in\mathbb{T}_{n}$. In particular, $\operatorname{\mathsf{dom}}^{w}[t]=\operatorname{\mathsf{dom}}^{w}(u)$ for any cluster monomial $u$ with $\operatorname{\mathsf{supp}}(u)=[{\bf x}_{t}]$.
@@
For a seed $({\bf x}_{t},B_{t})$ of $\mathcal{A}$, we denote by $G_{t}=G_{t}^{t_{0}}$ the $G$-matrix of the vertex $t$ with respect to the rooted vertex $t_{0}$ and by
$\mathbb{R}_{\geq 0}G_{t}\coloneqq\mathbb{R}_{\geq 0}{\bf g}_{1;t}+\cdots+\mathbb{R}_{\geq 0}{\bf g}_{n;t}\subseteq\mathbb{R}^{n}$
the cone in $\mathbb{R}^{n}$ spanned by the columns of $G_{t}=({\bf g}_{1;t},\ldots,{\bf g}_{n;t})$. We also denote by
$\mathbb{Q}_{>0}G_{t}$ $\coloneqq\mathbb{Q}_{>0}{\bf g}_{1;t}+\cdots+\mathbb{Q}_{>0}{\bf g}_{n;t}\subseteq\mathbb{Q}^{n},$
$\mathbb{Z}_{\geq 1}G_{t}$ $\coloneqq\mathbb{Z}_{\geq 1}{\bf g}_{1;t}+\cdots+\mathbb{Z}_{\geq 1}{\bf g}_{n;t}\subseteq\mathbb{Z}^{n}.$
@@
Let $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ be a green mutation in $\mathcal{A}$. Then there exist two vectors ${\bf v},{\bf v}^{\prime}\in\mathbb{Z}_{\geq 1}^{n}$ such that $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$.
@@
Without loss of generality, we can assume that $k=1$. We write $G_{t}=({\bf g}_{1},{\bf g}_{2},\ldots,{\bf g}_{n})$, where ${\bf g}_{i}\in\mathbb{Z}^{n}$. Since $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})=\mu_{1}({\bf x}_{t},B_{t})$, the $G$-matrix $G_{t^{\prime}}$ has the form $G_{t^{\prime}}=({\bf g}_{1}^{\prime},{\bf g}_{2},\ldots,{\bf g}_{n})$.
By Theorem 3.10 (i), we know that $\mathbb{R}_{\geq 0}G_{t}$ and $\mathbb{R}_{\geq 0}G_{t^{\prime}}$ are two cones of dimension $n$ in $\mathbb{R}^{n}$ and they have a common facet $\mathcal{C}$ (cone of dimension n1n-1) given by $\mathcal{C}=\mathbb{R}_{\geq 0}{\bf g}_{2}+\ldots+\mathbb{R}_{\geq 0}{\bf g}_{n}$.
Let $S=diag(s_{1},\ldots,s_{n})$ be the fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}$ and ${\bf c}_{1;t}$ the first column of the $C$-matrix $C_{t}=C_{t}^{t_{0}}$. By Theorem 3.10 (iii), we have $SC_{t}S^{-1}(G_{t})^{\rm T}=I_{n}$, equivalently, $(G_{t})^{\rm T}(SC_{t}S^{-1})=I_{n}$. So the first column vector $s_{1}^{-1}(S{\bf c}_{1;t})$ of $SC_{t}S^{-1}$ is an inner normal vector of the facet $\mathcal{C}$ of the cone $\mathbb{R}_{\geq 0}G_{t}$. So when we go along the direction $s_{1}^{-1}(S{\bf c}_{1;t})$, we can cross the common facet $\mathcal{C}$ from the cone $\mathbb{R}_{\geq 0}G_{t^{\prime}}$ to the cone $\mathbb{R}_{\geq 0}G_{t}$.
Now we take a vector ${\bf h}^{\prime}\in\mathbb{Z}_{\geq 1}G_{t^{\prime}}\subseteq\mathbb{R}_{\geq 0}G_{t^{\prime}}$. Then there exists $\lambda\in\mathbb{Q}_{>0}$ such that
${\bf h}\coloneqq{\bf h}^{\prime}+\lambda(s_{1}^{-1}(S{\bf c}_{1;t}))\in\mathbb{Q}_{>0}G_{t}\subseteq\mathbb{R}_{\geq 0}G_{t}.$
We can take a positive integer $p$ large enough such that $p{\bf h}\in\mathbb{Z}_{\geq 1}G_{t}$. Since $p{\bf h}\in\mathbb{Z}_{\geq 1}G_{t}\subseteq\mathbb{Z}^{n}$ and $p{\bf h}^{\prime}\in\mathbb{Z}_{\geq 1}G_{t^{\prime}}\subseteq\mathbb{Z}^{n}$, we have $p\lambda s_{1}^{-1}(S{\bf c}_{1;t})=p{\bf h}-p{\bf h}^{\prime}\in\mathbb{Z}^{n}$. Since $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ is a green mutation, we have ${\bf c}_{1;t}\in\mathbb{Z}_{\geq 0}^{n}$. So we have $p{\bf h}-p{\bf h}^{\prime}=p\lambda s_{1}^{-1}(S{\bf c}_{1;t})\in\mathbb{Z}_{\geq 0}^{n}$.
Since $p{\bf h}\in\mathbb{Z}_{\geq 1}G_{t}$ and $p{\bf h}^{\prime}\in\mathbb{Z}_{\geq 1}G_{t^{\prime}}$, there exist ${\bf v},{\bf v}^{\prime}\in\mathbb{Z}_{\geq 1}^{n}$ such that
$p{\bf h}=G_{t}{\bf v}\;\;\;\text{and}\;\;\;p{\bf h}^{\prime}=G_{t^{\prime}}{\bf v}^{\prime}.$
Thus $G_{t}{\bf v}-G_{t^{\prime}}{\bf v}^{\prime}=p{\bf h}-p{\bf h}^{\prime}\in\mathbb{Z}_{\geq 0}^{n}$. ∎